# A First Look at GPT Apps: Landscape and Vulnerability The original version was uploaded on Feb 23, 2024

### v2

@@ 3.2. Dataset
Once we capture web pages and responses from GPT apps, we parse the HTML pages and JSON files to extract critical information such as the app’s metadata, user feedback, and GPT app configurations. All the attributes and their possible values are detailed in Table 1. Metadata provides basic information about GPT apps, sourced from gptstore.ai. It typically includes details such as the app’s name, category, creator, features, etc. GPT app configurations refer to the data used by creators to build and refine their apps, often including proprietary algorithms and system prompts. User feedback, which includes metrics like rating score, rating ratio, and total rating counts, offers insights into user preferences and is primarily collected from the OpenAI Store. This comprehensive data collection and analysis allow us to understand the structure and dynamics within the LLM app ecosystem.
@@ 4. GPT Apps Statistics
4.1. Metadata
In this section, we analyze three types of data: metadata, GPT app configurations, and user feedback, uncovering their dynamics and takeaways.
Figure 3. Trend of Increasing Creators and Distribution of Creators that own different numbers of apps. The labels ’=1’, ’(1,10]’, and ’¿10’ denote creators with exactly one, between two and ten, and more than ten GPT apps, respectively.
4.1.1. Creator
Figure 6 shows that the growth in the number of creators within gptstore.ai starts off strong but eventually stabilizes. This increasing trend indicates a growing interest and engagement in developing GPT apps and suggests that more individuals and entities are participating in the GPT ecosystem, contributing to its diversity and expansion.
However, the distribution of GPT apps among these creators is markedly uneven. More than 97% of creators manage fewer than 10 GPT apps per person, with 68% and 29% creators managing 1 app and 2-10 apps separately, whereas only less than 5% of creators have more than 10 GPT apps. This distribution is even more skewed compared to that of Android apps in the Google Play Store (Wang et al., 2017), particularly with a substantial number of creators producing only one GPT app. Such the distribution exemplifies a ”long tail” scenario (Zhong and Michahelles, 2013), where many participants are involved at a small scale while a few dominate by volume.
$\bullet$ Findings: The rapid increase in creators during the initial phase highlights the strong enthusiasm for developing apps, also indicating the low barriers to entry in this market. Unlike mobile app development, creating GPT apps requires less specialized software development knowledge and shortens the time needed to build a new app, making it accessible to a broader audience. However, scaling up these projects can be challenging without significant resources or unique innovations.
4.1.2. Apps
Figure 4. Increasing Trend of GPT App Numbers in gptstore.ai. The labels ”Add”, ”Delete”, and ”Update” represent the number of GPT apps that have been newly created, removed, or modified, respectively, comparing to the last recording.
The growth trend of GPT apps, as depicted in Figure 4, parallels that of creators, beginning rapidly and then stabilizing. We observe that initially, there is a significant amount of activity, with many apps being added, updated, or deleted as the market adjusts to new market entries and demands. Over time, these activities decrease in frequency and scale, suggesting that the ecosystem reaches a level of maturity and requires fewer changes. However, achieving maturity within half a year is unusual in the dynamic field of technology and app development. Further analysis of app changes per creator reveals that initially, creators are highly enthusiastic, with some launching over 300 new apps per week. However, in later stages, the growth slows considerably, with only hundreds of apps added across the entire app store.
$\bullet$ Findings: The observation of fewer newly added and updated apps over time indicates a possible decline in creator enthusiasm for developing new GPT apps. This trend may be attributed to challenges in sustaining innovation or diminishing interest among creators. As the initial excitement of entering a new market wears off, the difficulties in continuously innovating and maintaining engagement can lead to a reduced pace of development and updates.
Takeaways: App stores should implement several strategies to maintain and boost creator engagement. For example, offering tangible incentives, such as financial rewards, grants, or profit-sharing models, can attract and retain creators by providing financial motivation. Hosting contests with attractive prizes can also ignite creativity and innovation, encouraging creators to experiment with new ideas and advance the capabilities of their apps. Additionally, providing opportunities for recognition—such as showcasing successful apps—can further motivate developers by celebrating their achievements and highlighting their contributions to the community. These initiatives not only foster a more vibrant and innovative marketplace but also help sustain the interest and involvement of app creators.
4.1.3. Conversation Counts
Figure 5. End-of-Month Distribution of Conversation Counts Among Top Kof Apps in gptstore.ai
Conversation counts, reflecting the number of daily interactions for each app, serve as a crucial indicator of app popularity. Figure 5 illustrates the distribution of accumulated conversation counts among Top $K\%$ apps from January through April, where $K\in[1,100]$. There is a sharp initial increase in conversation counts which then levels off, showing that a small fraction of top apps dominate the majority of user interactions. Specifically, the top 1% and 5% of apps are responsible for 90% and 95% of daily conversations, respectively. By the end of April, the Top 5 apps alone account for 25% of daily conversations and the Top 50 apps comprise 60%.
$\bullet$ Findings: The highly uneven distribution underscores a significant concentration of user engagement within a small group of apps, indicating that the vast majority of apps are not promoted well to the public.
Takeaways: To promote a more balanced and diverse app ecosystem, app stores could implement recommendation algorithms to better advertise apps to users, particularly new and less-known apps. Additionally, app stores could also highlight new apps through featured listings or promotional events, encouraging users to explore beyond the top few. This approach not only aids in increasing the visibility of new apps but also helps in maintaining a vibrant and dynamic app ecosystem.
4.1.4. Category
Figure 6. Distribution of GPT apps in Top N Categories within gptstore.ai
Figure 6 illustrates a consistent increase in the number of GPT apps across all categories within gptstore.ai. Furthermore, it highlights an uneven distribution among these categories, as shown by a concave-upward curve. Specifically, the top 10 categories encompass approximately 60% of all apps, and the top 20 categories comprise about 80%. This concentration of apps within a small subset of categories exemplifies the Pareto Principle (Sanders, 1987), which demonstrates that roughly 80% of consequences stem from 20% of causes. This pattern of app distribution is consistent with trends observed in other major platforms, such as Google (Carbunar and Potharaju, 2015). Notably, the last two categories in our dataset consistently contain no GPT apps, suggesting inefficiencies in how these categories are defined or utilized.
OpenAI lists the top 16 apps in each category, each boasting over 100,000 conversation counts. Due to the app store’s practice of consistently featuring these popular apps, an increasing number of users are drawn to try them, thereby further elevating their conversation counts. However, this approach restricts visibility for other apps within the same category, which complicates user navigation and limits exposure to a broader array of app options.
$\bullet$ Findings: The distribution of GPT apps within gptstore.ai is also concentrated. OpenAI’s strategy of consistently featuring the top 12 apps in each category may restrict visibility for other apps within the same categories, complicating user navigation and limiting exposure to a wider range of app options.
Features and Functions: GPT apps are equipped with five major functions and various customized APIs to enhance their performance. Here is a description of these five key functions. File Attachment: Users can upload files to apps. File Browser: This function enables users to access the web during conversations. DALL·E: Image generation capabilities are enabled with DALL·E. Python: Apps have the capability to write and execute Python codes. Knowledge File: GPT apps contain knowledge files that store essential information,
4.2. GPT App Configurations
@@
For the top 10,000 GPT apps, we conduct TriLevel configuration extraction to acquire system prompts, knowledge files, and external APIs. For the weekly top 500 GPT apps, our efforts are concentrated solely on system prompts, as information about external APIs and knowledge files is readily available in gptstore.ai.
4.2.1. Top 10,000
@@
Out of the 10,000 GPT apps we examined, 7706 are still publicly accessible. The remaining 2,294 apps may have been deleted or made private by their creators. While all of these apps include system prompts, only 34% (2,638 apps) have knowledge files, and just 2.6% (202 apps) use external APIs.
@@
Table 2. Success Rates of Configuration Extraction among the Top 10,000 GPT Apps. System prompts are sequentially extracted through a three-level prompt query.
Success Rate. As detailed in Table 2, we find that system prompts and file names are easy to access, whereas obtaining detailed file contents and API information proves more difficult.
In Level 1, we retrieve system prompts from over 87% of GPT apps. Progressing to Levels 2 and 3, we extract additional system prompts from 148 and 39 GPT apps, respectively. Ultimately, we successfully extract system prompts in nearly 90% of cases, leaving a 10% gap due to robust defenses.
For knowledge files, we successfully obtain 88% of the file names and download links. Nonetheless, the success rate for acquiring file contents is only 12.7% since the download process may encounter various unknown issues, hindered by the complexities of ChatGPT’s back-end architecture.
Regarding external APIs, we successfully obtain information from 81 out of 202 GPT apps, highlighting specific challenges in accessing full API capabilities.
System prompts. Figure 9 illustrates the distribution of prompt lengths among the top 10,000 GPT apps. It shows that over 76% of these apps feature prompts shorter than 2,000 words, indicating a preference for conciseness among developers. This trend suggests that the majority of GPT apps are designed to be straightforward, aiming to clearly convey their functionalities with minimal complexity.
@@
Figure 9. Popular Types of Knowledge Files.
Knowledge files. Knowledge files serve as supplementary resources, providing additional information that allows apps to generate more accurate, informed, or contextually relevant responses. Figure 9 shows that PDFs are the most common type of knowledge file, making up 42% of all extracted files with a total of 4,571 PDFs. Other frequently used formats are DOCX, JSON, CSV, and HTML, which are primarily text-based, rather than audio or video. Additionally, many GPT apps utilize multiple knowledge files to broaden their capabilities. For instance, Godot Assistant [^3] https://chat.openai.com/g/g-0e9nkbuRV-godot-assistant incorporates 44 files totaling over 23 million tokens.
4.2.2. Weekly Top 500
Figure 10. Success Rate for Acquiring System Prompts in the Top 10000 Apps and Weekly Top 500 Apps.
@@ 4.2.2. Weekly Top 500
Figure 11. Comparison of Common and Unique Apps in Weekly Top 500 Dataset. We compare apps from every two adjacent dates. Specifically, the data from March 6 is compared to the top 10,000 apps.
Common Apps. In the analysis of the top 500 apps conducted weekly over a period of nine weeks, 317 apps consistently appear in each dataset. System prompts are successfully extracted from 223 of these apps, while extraction fails for 44 apps. Configuration extraction consistently fails for 36 apps, whereas the remaining 14 apps initially succeed but fail in subsequent attempts.
Of the 223 GPT apps from which system prompts are consistently retrieved, 66 maintain identical prompts throughout the period. Additionally, 170 apps exhibit system prompts that are nearly identical, with a cosine distance greater than 0.995, indicating minimal variation. Meanwhile, 53 apps have updated their prompts over time, reflecting ongoing adjustments to their configurations.
$\bullet$ Findings: The high success rate in extracting configurations from GPT apps highlights their vulnerabilities, indicating that their protective mechanisms are insufficient. Among the components of GPT app configurations, system prompts are the easiest to access, followed by knowledge files, and then external APIs.
Takeaways: Some protective mechanisms should be implemented to enhance apps’ robustness against data extraction. For app stores, they can establish rigorous security protocols and review processes that evaluate apps’ susceptibility to attack before they are listed. For creators, incorporating protective instructions within system prompts can help bolster security. Additionally, employing external APIs to handle more sensitive data is another effective strategy to enhance protection.
@@ 4.3. User Feedback
Figure 12. Distribution of GPT Apps Across Four Different Ranges of Rating Counts: [1,10),[10,100),[100,500),500[1,10),[10,100),[100,500)$\geq$ 500[ 1 , 10 ) , [ 10 , 100 ) , [ 100 , 500 ) , ≥ 500.
User feedback is a critical indicator of both user preferences and the overall quality of an application. In March 2024, OpenAI enables users to rate apps, facilitating a quantitative assessment of satisfaction.
Rating Counts: Rating counts reflect the number of users who have rated the app on a scale from 1 to 5 stars. However, less than 18% of apps have actually received ratings. Additionally, the distribution of these ratings is highly uneven. As shown in Figure 12, more than 70% of the rated apps have fewer than 10 rating counts, while 17% receive between 10 and 100 ratings. This suggests that many apps suffer from low engagement or have a limited user base.
Rating Score: Figure 13 illustrates the distribution of apps across four ranges of average rating scores. It reveals that approximately 80% of rated apps have high scores (4 or higher). Moreover, 15%, 3%, 1% of apps have rating scores in $[3,4),[2,3),[1,2)$, respectively. This distribution indicates a high level of user satisfaction with the apps available in the market. However, relying solely on overall rating scores as an indicator of an app’s quality can be misleading. Because ratings from a small number of users may introduce bias, failing to accurately reflect the true quality of apps.
Figure 13. Distribution of GPT Apps Across Four Different Ranges of Rating Scores: $[1,2),[2,3),[3,4),[4,5]$.
This system does not currently include an option for users to provide reviews and comments. Detailed user reviews could significantly reveal the shortcomings of applications and drive enhancements.
$\bullet$ Findings: Users appear to be satisfied with apps in the market, as evidenced by the high proportion of rated apps receiving high scores. However, user feedback is concentrated on a relatively small number of apps, which could introduce bias into the ratings.
Takeaways: App stores can adopt strategies that increase user engagement with app ratings. For instance, they could offer incentives to encourage active participation in rating apps. Another approach is to use pop-up reminders after a user has spent some time with an app, prompting them to provide feedback. Most importantly, app stores should enable users to leave reviews and comments. This feedback is crucial for developers aiming to update their apps in response to user needs, ultimately enhancing the LLM app ecosystem.
@@ 4.4. Correlationship
From our analysis of metadata (§4.1), configuration (§4.2), and user feedback (§4.3), we observe a significant concentration of user engagement within a small number of apps. To further explore the relationships between various characteristics of GPT apps, we examine nine features: rating scores (Star), conversation counts (ConvCount), rating counts (RateCount), the use of external APIs (APIs), and the utilization of functions including python, browser, DALL·E image generation (Dalle), file attachments (FileAttach), knowledge files (KnowFile). We employed the Spearman Correlation (Myers and Sirois, 2004) to analyze their correlations. Below are our findings:
Figure 14. Spearman Correlation Matrix of GPT App Characteristics
1) There is a very strong correlation (0.98) between rating scores (Star) and rating counts (RateCount), suggesting that high-score apps tend to attract more feedback and user engagement.
2) App popularity (Conversation Counts) correlates positively with rating scores (Star) and rating counts (RateCount), with coefficients of 0.5 and 0.52, respectively. This suggests that more popular apps not only achieve higher rating scores but also garner more user feedback.
3) A strong correlation (0.90) between the use of APIs and the inclusion of knowledge files suggests that apps leveraging APIs are likely to also incorporate knowledge files. This may indicate that some APIs are employed to analyze or manage these files.
4) Moderate correlations between functions such as DALL·E, file attachments, Python, and Browser suggest that GPT apps are increasingly integrating multiple features to enhance app performance and user experience.
5. GPT Apps Plagiarism in the Wild

### v3

@@ 3.2. Dataset
Due to the large number of GPT apps, we focus on the most popular ones. We rank apps by conversation counts and extract configurations from the top 10,000 in February 2024. From March to August 2024, we conduct weekly extractions for the top 500 apps for longtitude analysis. In total, we collect data from the initial 10,000 apps and 21 weekly snapshots of the top 500 apps.
After capturing web pages and responses from GPT apps, we parse the HTML and JSON files to extract key information like metadata, user feedback, and app configurations. Table 1 displays all the attributes and corresponding descriptions. This dataset helps us analyze the structure and dynamics of the LLM app ecosystem.
4. GPT App Stores
Our research targets three GPT app stores: GPTStore.AI, GPTShunter, and OpenAI Store. The OpenAI Store lists 7 categories with the top 16 apps in each. GPTShunter also lists 7 categories but provides access to all apps. GPTStore.AI offers 50 categories with full access to all apps.
Figure 3. Number of Apps and Creators in Two GPT App Stores Over a 10-Month Monitoring Period.
4.1. Apps and Creators
Since the OpenAI Store does not provide a complete list of all available data, we monitor all available GPT apps and creators in GPTStore.AI and GPTShunter. Figure 3 compares the number of creators and apps in these two stores from January to October 2024. We observe that both GPTShunter and GPTStore.AI experience rapid growth in both creators and apps from January to May and then stabilize. This trend reflects a growing interest in developing GPT apps, with more individuals and organizations contributing to the ecosystem, which is gradually becoming more mature. However, GPTShunter has a significantly larger user base compared to GPTStore.AI, highlighting its greater popularity.
Additionally, we find that only 25% of apps receive updates, while the majority remain unchanged. This indicates that while there is initial enthusiasm for app creation, many apps become stagnant without further development or improvement. To maintain engagement and ensure continued growth, it would be beneficial for GPT app stores to provide incentives, such as financial rewards or feature enhancements, encouraging creators to update and improve their apps actively.
4.2. App Categories
GPTStore.AI lists 50 categories, detailed in Appendix B.1. Our analysis reveals that the top 10 categories account for about 55% of all apps, and the top 20 categories encompass 70%. This uneven distribution reflects the Pareto Principle (Sanders, 1987), where a small number of categories dominate the majority of app listings. Besides, this trend aligns with other major platforms like Google (Carbunar and Potharaju, 2015), where app distribution similarly follows a power-law pattern.
GPTShunter and OpenAI store list the same 7 categories: Writing, Productivity, Research&Analysis, Education, Lifestyle and Programming, yet the distribution remains highly unbalanced. In October, GPTShunter shows 82K+ apps in Education, 63K+ in Productivity, and 52K+ in Writing, 51K+ in Research while Programming (34K+), DALL·E (19K+) have significantly fewer apps. Additionally, over 160K apps remain uncategorized, making it difficult for users to discover them based on preferences.
$\bullet$ Findings: while GPTShunter and GPTStore.AI rapidly grow experience rapid growth, they now stabilize and only a small portion of apps receive updates, indicating a need for incentives to keep creators engaged. The uneven distribution of apps across 50 categories in GPTStore.AI and 160K uncategorized apps in GPTShunter highlights the need for better categorization and more active development to keep the ecosystem dynamic.
@@ 5. GPT Apps Statistics
For the top 10,000 GPT apps, we use TriLevel configuration extraction to obtain system prompts, knowledge files, and external APIs via OpenAI Store chatbot. For the weekly top 500 GPT apps, we focus only on system prompts, as API and knowledge file information are already available in GPTStore.AI and GPTShunter.
5.1. Configuration Extraction
5.1.1. Top 10,000 GPT Apps
@@
Out of the 10,000 GPT apps we examined, 7706 were still publicly accessible when retrieved in February 2024. The remaining apps may have been deleted or made private by their creators. While all of these apps include system prompts, only 34% (2,638 apps) have knowledge files, and just 2.6% (202 apps) use external APIs.
@@
Table 2. Success Rates of Configuration Extraction Among the Top 10,000 GPT Apps.
Success Rate. As detailed in Table 2, system prompts and file names are easy to access, but detailed file contents and API information are harder to obtain. We retrieve system prompts from 87% of apps in Level 1, with additional prompts from 148 and 39 apps in Levels 2 and 3, reaching nearly 90% overall. For knowledge files, we successfully obtain 88% of the file names and download links but only 12.7% of the actual file contents due to backend issues. For external APIs, we retrieve information from 81 out of 202 GPT apps, highlighting challenges in accessing full API capabilities.
System prompts. Figure 5 shows the distribution of prompt lengths among the top 10,000 GPT apps. Over 76% of these apps have prompts shorter than 2,000 words, indicating a preference for concise and simple design. This suggests most apps are built to clearly communicate their functionality with minimal complexity.
@@
Figure 5. Popular Types of Knowledge Files in Top 10,000 GPT apps.
Knowledge files. Knowledge files provide supplementary information, helping apps generate more accurate and relevant responses. Figure 5 shows that PDFs are the most common format, making up 42% of all extracted files (4,571 PDFs). Other common formats include DOCX, JSON, CSV, and HTML, which are primarily text-based. Besides, based on our discovery, many GPT apps use multiple knowledge files to enhance their capabilities.
5.1.2. Weekly Top 500 GPT Apps
Figure 6. Success Rate for Acquiring System Prompts in the Top 10,000 Apps and Weekly Top 500 Apps.
@@ 5.1.2. Weekly Top 500 GPT Apps
Figure 7. Comparison of Common and Unique Apps in Weekly Top 500 Dataset. Apps from each two adjacent dates are compared.
Common Apps. In May, after nine weeks of monitoring, 317 out of 500 apps consistently appear in every dataset. We successfully extract system prompts from 229 apps all the time, while extraction consistently fails for 44 apps. 10 apps allow prompt extraction initially but fail later, and 57 apps update their prompts over time.
In August, after 21 weeks of monitoring, 196 out of 500 apps consistently appear in every dataset. We successfully extract system prompts from 110 apps all the time, while extraction fails consistently for 12 apps. Additionally, 36 apps initially allow prompt extraction but fail later and 36 apps update their prompts over time.
Our analysis shows that many GPT apps consistently appear over time, indicating their stability or popularity. Some apps initially allow system prompt extraction but later block access, suggesting evolving security practices by developers to protect their configurations. Additionally, frequent prompt updates indicate that developers are actively refining functionality to meet user needs, showing a dynamic and responsive ecosystem where apps continuously adapt to improve performance and maintain relevance.
$\bullet$ Findings: The high success rate in extracting GPT app configurations reveals vulnerabilities and weak protective mechanisms. System prompts are the easiest to access, followed by knowledge files and external APIs. To improve security, app stores should implement stricter review processes, and creators should incorporate protective instructions in prompts or use external APIs for sensitive data.
5.2. Conversation Counts
Conversation counts, reflecting the daily interactions for each app, serve as a crucial indicator of app popularity.
Figure 8. Distribution of GPT Apps Across Four Different Ranges of Conversation Counts: ¡=100, [100,1K), [1K,10K), ¿=10K.
Figure 8 compares the distribution of GPT apps based on their conversation counts and highlights a skewed distribution of popularity in two stores. In both stores, most (over 80%) apps have fewer than 100 conversation counts (gray), indicating low user engagement. A small number of apps fall into the higher conversation count ranges, with a slight increase in apps receiving between 1,000 and 10,000 conversations (green) and very few exceeding 10,000 conversations (blue).
$\bullet$ Findings: The GPT app ecosystem is dominated by a few highly popular apps, while most apps receive little user engagement. This suggests a need for strategies to improve visibility and user interaction for less popular apps to create a more balanced ecosystem.
@@ 5.3. User Feedback
Figure 9. Distribution of GPT Apps Across Four Different Ranges of Rating Counts: $10),[10,100),[100,500),\geq 500[ 1 , 1$0 ) , [ 10 , 100 ) , [ 100 , 500 ) , ≥ 500.
Figure 10. Distribution of GPT Apps Across Four Different Ranges of Rating Scores: [1,2),[2,3),[3,4),[4,5]12233445[1,2),[2,3),[3,4),[4,5][ 1 , 2 ) , [ 2 , 3 ) , [ 3 , 4 ) , [ 4 , 5 ].
User feedback is a key indicator of user preferences and app quality. In March 2024, OpenAI introduced a 5-star rating system, allowing users to provide quantitative assessments of their satisfaction with GPT apps.
Rating Score. Figure 10 compares user ratings distribution in two stores. Both show that over 80% of apps (gray) remain unrated. Most rated apps fall in the [4, 5) range, showing high user satisfaction on both platforms. Ratings in the [3, 4) and [2, 3) ranges are fewer, with very few in the [1, 2) range. The similar patterns across platforms highlight the need for more user engagement due to the large number of unrated apps.
Rating Counts. Figure 9 compares the rating count distribution in two stores. Over 80% of apps (gray) on both platforms have no ratings. Among rated apps, most fall into [1, 10) (red) range, indicating even rated apps receive few reviews. A smaller portion of apps has between [10, 100) (blue) and [100, 500) (green) ratings, with very few exceeding 500 (pink). The trend is similar across both stores, showing minimal user engagement for most apps.
$\bullet$ Findings: In app stores, user engagement in feedback is low, with over 80% of apps unrated and most rated apps receiving few reviews. This lack of participation limits app discoverability and developer feedback, highlighting the need for strategies to boost user ratings and reviews.
@@ 5.4. Correlationship
Our analysis of metadata (§4), configuration (§5.1), and user feedback (§5.3) reveals a concentration of user engagement within a small number of apps. To further explore relationships between different app characteristics, we examine nine features: rating scores (Star), conversation counts (ConvCnt), rating counts (RateCnt), external APIs (APIs), and functions like Python, browser, DALL·E (Dalle), file attachments (FileAtt), and knowledge files (KnowFile). Using Spearman Correlation, we analyze how these features relate to each other. Below are our findings:
Figure 11. Correlation Matrix of App Configurations.
1. There is a strong correlation (0.96) between rating scores (Star) and rating counts (RateCnt), indicating that high-rated apps tend to receive more user feedback.
2. App popularity (ConvCnt) correlates positively with rating scores (0.56) and rating counts (0.62), suggesting that popular apps receive higher ratings and more feedback.
3. A strong correlation (0.89) between APIs and knowledge files suggests that apps using APIs often incorporate knowledge files, possibly for managing or analyzing them.
4. Moderate correlations between features like DALL·E, file attachments, Python, and Browser indicate that GPT apps are integrating multiple functions to improve performance and user experience.
6. Plagiarism in the Wild